OpenStudy (anonymous):

Is (2, 0) a solution to y = -x + 2?

ganeshie8 (ganeshie8):

plugin x=2, y=0 if the equation satisfies then its a solution

OpenStudy (anonymous):

Calc. You messed it up.

OpenStudy (anonymous):

It's y = -x + 2 therefore should be 0 = -2 + 2 I don't see how you got a +2 times +2

OpenStudy (anonymous):

Yeah. I see. Plug in the coordinate (2, 0). If it is true, then it is a solution. \[y = -x + 2\]\[0 = -(2) + 2\]Is this statement true?